Grade Upto college level Algebra

Find the number of ways of selecting a cricket team of 11 players from 7 batsmen and 6 bowlers such that there will be at least 5 bowlers in the team.

Manvendra Singh chahar
12 Years agoGrade Upto college level

1 Answer

SHAIK AASIF AHAMED
12 Years ago
Hello student,
Since the team consists of at least 5 bowlers, the selection may be of the following types.
Batsmen (7) Bowlers (6)
Type 1 6 5
Type 2 5 6
The number of selections in first type =6c5* 7c6 =42
The number of selections in second type=7c5*6c6=21
∴The required number of ways selecting the cricket team = 42 + 21 = 63.
Thanks and Regards
Shaik Aasif
askIITians faculty